Using Winkler method, oxygen fixed sample is to be tested for dissolved oxygen in the laboratory.
If 12.5 mL of a 0.02N sodium thiosulfate titrant is used for 200 mL of sample, what is the dissolved
oxygen content of this sample?
1
Expert's answer
2020-10-16T08:59:34-0400

1 mole of O2 is equivalent to 4 moles of sodium thiosulfate.

Since the gram equivalent weight of sodium thiosulfate is 1 the molarity of 0.02N solution will be 0.02M


So to find the moles of thiosulfate,

=(12.5÷1000)0.02

=0.000025 moles.

The moles for oxygen will be obt

(0.000025÷4)×1 =0.00000625 moles

This are present in 200ml of the sample.

To find the amount for 1L

(1000÷200)×0.00000625= 0.00003125moles

And to express this moles in grams, it's multiplied by the molar mass of oxygen

0.00003125×32=0.001g

Therefore the dissolved oxygen is 1mg/L